Showing all work, determine whether the lines given by the equations below are parallel, perpendicular, or neither:
y = 2/3x - 6
6x + 4y = 5

Answer by unlockmath(1688)

Hello,
Best way is to get each equation in Slope-intercept form.
y = 2/3x - 6
6x + 4y = 5
First equation is already.
This one (6x + 4y = 5), subtract 6x and divide by 4 to get:
y=-2/3y+5
Answer: neither
Make sense?
RJ
www.math-unlock.com